OpenStudy (anonymous):

which property is represented? 0+45=45

jimthompson5910 (jim_thompson5910):

This is the additive identity property

OpenStudy (anonymous):

huh? never heard of it. i'm tired wish i didn't have homework. you sure?

jimthompson5910 (jim_thompson5910):

are you given a list of choices?

OpenStudy (anonymous):

nope. that would be a lot easier

jimthompson5910 (jim_thompson5910):

yes I'm sure, but your book may call it something else

OpenStudy (anonymous):

oh. well i don't want to get it wrong

jimthompson5910 (jim_thompson5910):

The additive identity property states that 0 + x = x Where x is any number

jimthompson5910 (jim_thompson5910):

and 0 is known as the additive identity

jimthompson5910 (jim_thompson5910):

basically, adding zero to a number does not change that number

OpenStudy (anonymous):

oh you are really smart I don't like math

jimthompson5910 (jim_thompson5910):

well I'm sure with more practice you'll get good at it too

OpenStudy (anonymous):

people say I am I just get distracted really easy. most of the time I don't ask so many questions

jimthompson5910 (jim_thompson5910):

the most important thing to remember is that you just need to get things translated into terms you can understand

jimthompson5910 (jim_thompson5910):

learning math is like learning a new language

OpenStudy (anonymous):

Yep. while we are talking about it is it somthing like the equivilant property or somthing for this question (r+s)+t=r+(s+t)

jimthompson5910 (jim_thompson5910):

(r+s)+t=r+(s+t) is using the idea that parenthesis can be shifted around when adding (ie you can add numbers in any order and it doesn't matter) This is known as the associative property of addition

OpenStudy (anonymous):

Oh yeah!!!!! I remember now!

jimthompson5910 (jim_thompson5910):

awesome

OpenStudy (anonymous):

I got confused with this last year too

jimthompson5910 (jim_thompson5910):

i guess you can think of parenthesis "associating" (ie pairing/grouping them) the terms r and s on the left side then associating s and t on the right side

OpenStudy (anonymous):

this is associative too isn't it 84+3=3+84 +=times

jimthompson5910 (jim_thompson5910):

this is the commutative property of multiplication

jimthompson5910 (jim_thompson5910):

btw use * to mean multiplication

OpenStudy (anonymous):

ok I couldn't find it I knew that somtimes I have blonde moments

jimthompson5910 (jim_thompson5910):

shift + 8 or the * key on the numlock pad

OpenStudy (anonymous):

Yep. thanks I may have another question so stay online :)

jimthompson5910 (jim_thompson5910):

ok